### v3

@@ 1.4 Paper Organization
We introduce the preliminaries in Section 2, including popular measures of the calibration error, decision making and swap regret, and the online binary calibration problem. We introduce Maximum Calibration Loss (CDL) in Section 3 and discuss its alternative formulation using Bregman divergences and its approximation via V-shaped Bregman divergences, which will be useful to establish our main result. Section 4 discusses the connection between CDL and other calibration error metrics. In Section 5, we present our main result, an efficient online binary prediction algorithm that guarantees $O(\frac{\log T}{\sqrt{T}})$ CDL. The key technical idea behind this result is a lemma (Lemma 5.2) we prove in Section 5.1 which allows us to attribute CDL to bucket-wise biases. Additionally, we give a non-constructive but simpler minimax proof of the $O(\frac{\log T}{\sqrt{T}})$ CDL guarantee in Appendix A. This simpler proof also crucially relies on our key technical lemma (Lemma 5.2) in Section 5.1.
@@ 1.5.2 Online Calibration Algorithms
Recent research has made significant progress in proving upper and lower bounds on the optimal rate achievable for both ECE and the distance to calibration in online binary prediction, though significant gaps remain between the current best upper and lower bounds. For ECE minimization, Foster and Vohra (1998) design a randomized algorithm that achieves $O(T^{-2/3})$ expected ECE, which remains the best known upper bound. Existence proofs and constructions of such algorithms have been further explored in several subsequent works (Hart, 2022; Foster and Hart, 2021). A recent work (Qiao and Valiant, 2021) show a lower bound of $\Omega(\frac{T^{0.528}}{T})$ to online ECE minimization. For the distance to calibration, Qiao and Zheng (2024) give a non-constructive minimax-based proof for an $O(\frac{1}{\sqrt{T}})$ upper bound and an $\Omega(T^{-1/3})$ lower bound for the same problem. Soon afterwards, Arunachaleswaran et al. (2024) provide a construction of an efficient algorithm that achieves the $O(\frac{1}{\sqrt{T}})$ upper bound for the distance to calibration.
The literature on online regret minimization for all downstream decision makers is technically closest to our paper. Kleinberg et al. (2023) observe that the swap regret of actions for any payoff-bounded decision task is linearly upperbounded by ECE. This observation quantitatively justifies the qualitative equivalence between no swap regret and calibration in Foster and Vohra (1997). However, the $\Omega(\frac{T^{0.528}}{T})$ lowerbound presents a barrier in efficient swap regret minimization via ECE minimization. To overcome this barrier from ECE, several relaxations of CDL have been considered by recent works to achieve $\tilde{O}(\frac{1}{\sqrt{T}})$ regret rates in online binary prediction. Kleinberg et al. (2023) achieve the $O(\frac{1}{\sqrt{T}})$ external regret maximized over payoff-bounded tasks, a necessary but insufficient condition for asymptotic calibration. On the other hand, Roth and Shi (2024), an independent work of ours, show $\widetilde{O}(|A|\frac{1}{\sqrt{T}})$ swap regret bounds that depend additionally on the number of actions $|A|$ in the downstream decision task, which can be loose when a decision task has many or even infinitely many actions.
@@ 2 Preliminaries
Throughout the paper, we use $p\in[0,1]$ to denote a prediction, and use $\theta$ to denote a binary state. A prediction $p$ can be viewed as a distribution over the state space $\Theta=\{0,1\}$, and we write $\theta\sim p$ when we sample $\theta$ from the Bernoulli distribution with mean $p$, i.e., $\Pr[\theta=1]=p$. For a real number $x$, we use $(x)_{+}$ or $[x]_{+}$ to denote $\max\{x,0\}$. We use $\mathbb{I}\left[\cdot\right]$ to denote the 0-1 indicator function: $\mathbb{I}\left[\text{statement}\right]=1$ if the statement is true, and $\mathbb{I}\left[\text{statement}\right]=0$ if the statement is false.
@@ 2.1 Measures of Calibration Error
$\textsc{ECE}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{q_{i}\in Q}n_{i}|q_{i}-\widehat{q}_{i}|.$
@@
Definition 2.5 (Smooth Calibration Error (Kakade and Foster, 2008)).
@@
Definition 2.6 (Distance to Calibration (Błasiok et al., 2023a)).
@@ 2.4 Calibration Fixed Decision Loss
$\textsc{PSwap}_{S}(\bm{p},\bm{\theta})=\max_{\sigma:\Delta(\Theta)\to A}\sum_{t\in[T]}\bigg{[}S(\sigma(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]},$
@@
$\textsc{Swap}_{U}(\bm{a},\bm{\theta})\leq\textsc{PSwap}_{S_{U}}$
@@
$\textsc{CFDL}_{S}=\textsc{PSwap}_{S}(\bm{p},\bm{\theta})\leq 2\textsc{ECE}(\bm{p},\bm{\theta}).$
@@ 2.5 Online Binary Calibration
We allow the predictor to be randomized, in which case we can view its strategy as a distribution ${\mathcal{F}}$ over deterministic strategies $F$. When we use an error metric, say MSR, to evaluate the predictions made by our predictor in $T$ rounds, our goal is to minimize the expected value of the error metric w.r.t. the worst-case adversary $A$, i.e., we want ${\mathbf{E}}_{F\sim{\mathcal{F}}}\left[\textsc{CDL}(h_{T}(F,A))\right]$ to be small for every adversary $A$.
@@
Theorem 2.21 (Foster and Vohra (1998), see also Hart (2022); Foster and Hart (2021)).
@@
Theorem 2.22 (Qiao and Valiant (2021)).
@@
Theorem 2.24 (Qiao and Zheng (2024); Arunachaleswaran et al. (2024)).
@@
Theorem 2.25 (Qiao and Zheng (2024)).
@@ 3.1 Bregman Divergence Formulation
Each proper scoring rule induces a Bregman divergence, measuring the loss from incorrect prediction. The Bregman divergence is defined with the convex utility function of a scoring rule. See Figure 2 for a geometric demonstration of proper scoring rules.
Theorem 3.4 (McCarthy (1956); Savage (1971)).
A scoring rule is proper if and only if there exists a convex utility function $u:\Delta(\Theta)\to\mathbb{R}$ such that
@@
Specified by a convex function $u:\Delta(\Theta)\to\mathbb{R}$, the Bregman divergence $\textsc{Breg}(p^{\prime},p)$ is defined as [^8] The definition of Bregman divergence here reverses the order of input in the conventional definition to align with proper scoring rules. In the conventional definition, $\textsc{Breg}(p,\widehat{p})=u(p)-u(\widehat{p})+\nabla u(\widehat{p})(\widehat{p}-p)$.
@@ 3.2 Approximation via V-Bregman Divergences
Figure 3: The thick black line plots the special convex utility function $u$ for the scoring rule. The convex utility function is V-shaped, consisting of two linear pieces intersecting at $\mu$. Once fixing the prediction $p$, the score $S(p,\theta)=u(p)+\nabla u(p)\cdot(\theta-p)$ is linear in the state $\theta$. The scoring rule offers two set of scores for the prediction to selecttwo linear lines for prediction. When $p\leq q_{0}$, the prediction selects scores $\{S(0,0),S(0,1)\}$. Otherwise, the prediction selects $\{S(1,0),S(1,1)\}$.
@@
Now let us consider any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$. By Theorem 3.4, there exists a convex function $u:[0,1]\to[0,1]$ with bounded sub-gradients $\nabla u(p)\in[-1,1]$ such that (11) holds. We can construct a sequence of twice-differentiable convex functions $u_{1},u_{2},\ldots$ such that for every $p\in[0,1]$, $u_{j}(p)\to u(p)$ and $u_{j}^{\prime}(p)\to\nabla u(p)$ as $j\to+\infty$. Let $S_{j}$ denote the proper scoring rule generated by $u_{j}$ as in (11). For every $p,\widehat{p}\in[0,1]$, we have
$\lim_{j\to\infty}\textsc{Breg}_{S_{j}}(p,\widehat{p})=\textsc{Breg}_{S}(p,\widehat{p}).$
Let $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ be a finite prediction space such that $p_{t}\in Q$ for every $t\in[T]$. Define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2). By Lemma 3.8,
@@
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0)).$
The right hand side converges as $j\to\infty$:
$\lim_{j\to\infty}\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))=\textsc{VCFDL}(\bm{p},\bm{\theta})(\nabla u(1)-\nabla u(0))\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
@@
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\sum_{q_{i}\in Q}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i})$ $=\lim_{j\to\infty}\sum_{q_{i}\in Q}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$
$\leq\lim_{j\to\infty}\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))$
@@ 3.3 Computation of CDL
$\max_{S}\qquad$ $\sum_{q\in Q,\theta\in\Theta}(s_{\widehat{q},\theta}-s_{q,\theta})\widehat{q}(\theta)$
@@ 5.1 Attributing CDL to Bucket-wise Biases
$\textsc{CDL}(\bm{p},\bm{\theta})\leq 4\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T}).$
@@
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq 2\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$ (19)
@@
$\leq 2\frac{m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$
@@ 5.2 Efficient CDL Minimization Algorithm
For each expert $(i,\sigma)$, the guarantee (26) is stronger than more standard guarantees for the experts problem in that the right hand side of (26) has an $\sqrt{n_{i}}$ dependence rather than an $\sqrt{T}$ dependence. For the auxiliary expert, we get the guarantee (25), which can be viewed as a special form of (26) with $l_{t,i,\sigma}=0$ and $n_{i}=0$.
Algorithm 1 Algorithm for MSR minimization.
@@ 7. Calculate expert gains $l_{t,i,\sigma}$ for every $i\in[m]$ and $\sigma=\pm 1$:
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]$ $\leq\frac{4m}{T}{\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T})$ (by Lemma 5.2)
$\leq O(\frac{m\log(mT)}{T})+\frac{4\alpha}{\sqrt{T}}+4\beta+O(\frac{\alpha^{2}m\log m}{T})$ (by Lemma 5.5)
$=O(\frac{m\log T}{T})+O(\frac{\sqrt{\log T}}{\sqrt{T}})+O(\frac{\log T}{\sqrt{T}})+O(\frac{\log T}{\sqrt{T}})$
$=O(\frac{\log T}{\sqrt{T}}).$

### v4

@@ 1.4 Paper Organization
We introduce the preliminaries in Section 2, including popular measures of the calibration error, decision making and swap regret, and the online binary calibration problem. We introduce Calibration Decision Loss (CDL) in Section 3. We discuss its alternative formulation using Bregman divergences and its approximation via V-shaped Bregman divergences, which will be useful to establish our main result. Section 4 discusses the connection between CDL and other calibration error metrics. In Section 5, we present our main result, an efficient online binary prediction algorithm that guarantees $O(\frac{\log T}{\sqrt{T}})$ CDL. The key technical idea behind this result is a lemma (Lemma 5.2) we prove in Section 5.1 which allows us to attribute CDL to bucket-wise biases. Additionally, we give a non-constructive but simpler minimax proof of the $O(\frac{\log T}{\sqrt{T}})$ CDL guarantee in Appendix A. This simpler proof also crucially relies on our key technical lemma (Lemma 5.2) in Section 5.1.
@@ 1.5.2 Online Calibration Algorithms
Recent research has made significant progress in proving upper and lower bounds on the optimal rate achievable for both ECE and the distance to calibration in online binary prediction, though significant gaps remain between the current best upper and lower bounds. For ECE minimization, Foster and Vohra (1998) design a randomized algorithm that achieves $O(T^{-2/3})$ expected ECE, which remains the best known upper bound. Existence proofs and constructions of such algorithms have been further explored in several subsequent works (Hart, 2022; Foster and Hart, 2021). A recent work (Qiao and Valiant, 2021) show a lower bound of $\Omega(T^{-0.472})$ to online ECE minimization. For the distance to calibration, Qiao and Zheng (2024) give a non-constructive minimax-based proof for an $O(\frac{1}{\sqrt{T}})$ upper bound and an $\Omega(T^{-1/3})$ lower bound for the same problem. Soon afterwards, Arunachaleswaran et al. (2024) provide a construction of an efficient algorithm that achieves the $O(\frac{1}{\sqrt{T}})$ upper bound for the distance to calibration.
The literature on online regret minimization for all downstream decision makers is technically closest to our paper. Kleinberg et al. (2023) observe that the swap regret of actions for any payoff-bounded decision task is linearly upperbounded by ECE. This observation quantitatively justifies the qualitative equivalence between no swap regret and calibration in Foster and Vohra (1997). However, the $\Omega(T^{-0.472})$ lowerbound presents a barrier in efficient swap regret minimization via ECE minimization. To overcome this barrier from ECE, several relaxations of CDL have been considered by recent works to achieve $\tilde{O}(\frac{1}{\sqrt{T}})$ regret rates in online binary prediction. Kleinberg et al. (2023) achieve the $O(\frac{1}{\sqrt{T}})$ external regret maximized over payoff-bounded tasks, a necessary but insufficient condition for asymptotic calibration. On the other hand, Roth and Shi (2024) show $O(|A|\sqrt{\frac{\log T}{T}})$ swap regret bounds that depend additionally on the number of actions $|A|$ in the downstream decision task, which can be loose when a decision task has many or even infinitely many actions.
@@ 2 Preliminaries
Throughout the paper, we denote a prediction by $p\in[0,1]$, and a binary state by $\theta\in\{0,1\}$. A prediction $p$ can be viewed as a distribution over the state space $\Theta=\{0,1\}$, and we write $\theta\sim p$ when we sample $\theta$ from the Bernoulli distribution with mean $p$, i.e., $\Pr[\theta=1]=p$. For a real number $x$, we use $(x)_{+}$ or $[x]_{+}$ to denote $\max\{x,0\}$. We use $\mathbb{I}\left[\cdot\right]$ to denote the 0-1 indicator function: $\mathbb{I}\left[\text{statement}\right]=1$ if the statement is true, and $\mathbb{I}\left[\text{statement}\right]=0$ if the statement is false.
@@ 2.1 Measures of Calibration Error
$\textsc{ECE}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{t\in[T]}|p_{t}-\sigma^{*}(p_{t})|.$
Equivalently, $\textsc{ECE}=\frac{1}{T}\sum_{q_{i}\in Q}n_{i}|q_{i}-\widehat{q}_{i}|$.
@@
Definition 2.5 (Smooth Calibration Error, Kakade and Foster, 2008).
@@
Definition 2.6 (Distance to Calibration, Błasiok et al., 2023a).
@@ 2.4 Calibration Fixed Decision Loss
$\textsc{PSwap}_{S}(\bm{p},\bm{\theta})=\max_{\sigma:\Delta(\Theta)\to A}\frac{1}{T}\sum_{t\in[T]}\bigg{[}S(\sigma(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]},$
@@
$\textsc{Swap}_{U}(\bm{a},\bm{\theta})\leq\textsc{PSwap}_{S_{U}}(\bm{p},\bm{\theta})$
@@
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\textsc{PSwap}_{S}(\bm{p},\bm{\theta})\leq 2\textsc{ECE}(\bm{p},\bm{\theta}).$
@@ 2.5 Online Binary Calibration
We allow the predictor to be randomized, in which case we can view its strategy as a distribution ${\mathcal{F}}$ over deterministic strategies $F$. When we use an error metric, say ECE, to evaluate the predictions made by our predictor in $T$ rounds, our goal is to minimize the expected value of the error metric w.r.t. the worst-case adversary $A$, i.e., we want ${\mathbf{E}}_{F\sim{\mathcal{F}}}\left[\textsc{CDL}(h_{T}(F,A))\right]$ to be small for every adversary $A$.
@@
Theorem 2.21 (Foster and Vohra, 1998, see also Hart, 2022; Foster and Hart, 2021).
@@
Theorem 2.22 (Qiao and Valiant, 2021).
@@
Theorem 2.24 (Qiao and Zheng, 2024; Arunachaleswaran et al., 2024).
@@
Theorem 2.25 (Qiao and Zheng, 2024).
@@ 3.1 Bregman Divergence Formulation
Each proper scoring rule induces a Bregman divergence, measuring the loss from incorrect prediction. The Bregman divergence is defined with the convex function of a scoring rule. See Figure 2 for a geometric demonstration of proper scoring rules.
Theorem 3.4 (McCarthy, 1956; Savage, 1971).
A scoring rule is proper if and only if there exists a convex function $u:[0,1]\to\mathbb{R}$ and its sub-gradient $\nabla u:[0,1]\to\mathbb{R}$ such that
@@
Specified by a convex function $u:[0,1]\to\mathbb{R}$ and its sub-gradient $\nabla u:[0,1]\to\mathbb{R}$, the Bregman divergence $\textsc{Breg}(p^{\prime},p)$ is defined as [^9] The definition of Bregman divergence here reverses the order of input in the conventional definition to align with proper scoring rules. In the conventional definition, $\textsc{Breg}(p,\widehat{p})=u(p)-u(\widehat{p})+\nabla u(\widehat{p})(\widehat{p}-p)$.
@@ 3.2 Approximation via V-Bregman Divergences
Equivalently, this is the Bregman divergence for the convex function $u:[0,1]\to[0,1]$ shown in Figure 3, whose sub-gradient is
$\nabla u(p)=\begin{cases}-\frac{1}{2\max\{1-\mu,\mu\}},&\text{if }p<\mu;\\ \frac{1}{2\max\{1-\mu,\mu\}},&\text{if }p\geq\mu.\end{cases}$
@@
Figure 3: The thick black line plots the special convex function $u$ for the scoring rule. The convex utility function is V-shaped, consisting of two linear pieces intersecting at $\mu$. Once fixing the prediction $p$, the score $S(p,\theta)=u(p)+\nabla u(p)\cdot(\theta-p)$ is linear in the state $\theta$. The scoring rule offers two set of scores for the prediction to selecttwo linear lines for prediction. When $p\leq q_{0}$, the prediction selects scores $\{S(0,0),S(0,1)\}$. Otherwise, the prediction selects $\{S(1,0),S(1,1)\}$.
@@
Let $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ be a finite prediction space such that $p_{t}\in Q$ for every $t\in[T]$. For $i=1,\ldots,m$, define $n_{i}$ and $\widehat{q}_{i}$ as in (1) and (2).
Let us consider any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$. By Theorem 3.4, there exists a convex function $u:[0,1]\to[0,1]$ with bounded sub-gradients $\nabla u(p)\in[-1,1]$ such that (11) holds. We can construct a sequence of twice-differentiable convex functions $u_{1},u_{2},\ldots$ to approximate $u$ in the following manner. For every $q\in\{0,1,q_{1},\ldots,q_{m},\widehat{q}_{1},\ldots,\widehat{q}_{m}\}$, the derivative $u_{j}^{\prime}(q)$ satisfies $u_{j}^{\prime}(q)=\nabla u(q)$ for every $j$, and the function value $u_{j}$ itself satisfies $u_{j}(q)\to u(q)$ as $j\to+\infty$. Let $S_{j}$ denote the proper scoring rule generated by $u_{j}$ as in (11). For every $i=1,\ldots,m$, we have
$\lim_{j\to\infty}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})=\textsc{Breg}_{S}(q_{i},\widehat{q}_{i}).$
By Lemma 3.8, for every $j=1,2,\ldots,$
@@
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(\nabla u(1)-\nabla u(0))$
$\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
@@
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i})$ $=\lim_{j\to\infty}\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$
@@ 3.3 Computation of CDL
$\max_{S}\qquad$ $\frac{1}{T}\sum_{q\in Q,\theta\in\Theta}(s_{\widehat{q},\theta}-s_{q,\theta})\widehat{q}(\theta)$
@@ 5.1 Attributing CDL to Bucket-wise Biases
$\textsc{CDL}(\bm{p},\bm{\theta})\leq\frac{4m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{4\alpha}{\sqrt{T}}+4\beta+O\Big{(}\frac{\alpha^{2}m\log m}{T}\Big{)}.$
@@
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$ (19)
@@
$\leq\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$
@@ 5.2 Efficient CDL Minimization Algorithm
For each expert $(i,\sigma)$, the guarantee (26) is stronger than more standard guarantees for the experts problem in that the right hand side of (26) has a $\sqrt{n_{i}}$ dependence rather than a $\sqrt{T}$ dependence. For the auxiliary expert, we get the guarantee (25), which can be viewed as a special form of (26) with $l_{t,i,\sigma}=0$ and $n_{i}=0$.
Algorithm 1 Algorithm for CDL minimization.
@@ 7. Calculate expert gains $l_{t,i,\sigma}$ for every $i\in[m]$ and $\sigma=\pm 1$:
${\mathbf{E}}\left[\textsc{CDL}(\bm{p},\bm{\theta})\right]$ $\leq\frac{4m}{T}{\mathbf{E}}\left[\mathcal{D}(\bm{p},\bm{\theta})\right]+\frac{4\alpha}{\sqrt{T}}+4\beta+O\Big{(}\frac{\alpha^{2}m\log m}{T}\Big{)}$ (by Lemma 5.2)
$\leq O\Big{(}\frac{m\log(mT)}{T}\Big{)}+\frac{4\alpha}{\sqrt{T}}+4\beta+O\Big{(}\frac{\alpha^{2}m\log m}{T}\Big{)}$ (by Lemma 5.5)
$=O\Big{(}\frac{m\log T}{T}\Big{)}+O\Big{(}\frac{\sqrt{\log T}}{\sqrt{T}}\Big{)}+O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}+O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}$
$=O\Big{(}\frac{\log T}{\sqrt{T}}\Big{)}.$